if sine theta = -4/7 and is in quadrant 3, then the sum of sine theta + sine (theta-pie)+sine(2pie+theta) is?
Answer by lwsshak3(11628)

if sine theta = -4/7 and is in quadrant 3, then the sum of sine theta + sine (theta-pie)+sine(2pie+theta) is?
***
sinx=-4/7 (Q3)
cosx=-√(1-(sin^2(x)=-√(1-16/49)=-√(33/49)=-√33/7
sinx+sin(x-π)+sin(2π+x)
sinx+(sinxcosπ-cosxsinπ)+(sin2πcosx+cos2πsinx)
=-4/7+(-4/7*-1-(-√33/7)*0)+(0*-√33/7+1*-4/7)
=-4/7+(4/7)-4/7
=-4/7